The Frequesy of a wave is 100Hz and its wavelength is 50m. Its velocity is

Grade:8

1 Answers

Raghav Rao
7 years ago
V=fλ  f=100Hz , λ=50m.
V=100*50 =5000 m/s.
[ we get this equation by the following way:-
V=dx/dt 
At t=T(time period) , x=λ .
So v=λ/T 
1/T=f
So, v=fλ ]